OpenStudy (anonymous):

find (g o f)(3),f(x)=3x-2, g(x)= x^2

OpenStudy (kittykitty04):

g(f(x)) -> g(f(3)) = (3(3)-2)^2 (3(3) -2) = 7, 7^2 = 49